A fighter jet is catapulted off an aircraft carrier from rest to 75 m/s. If the aircraft carrier deck is 100 m long, what is the acceleration of the jet?
1
Expert's answer
2020-10-23T11:50:56-0400

We can find the acceleration of the fighter jet from the kinematic equation:


"v^2=v_0^2+2as,"

here, "v_0=0\ \dfrac{m}{s}" is initial velocity of the fighter jet, "v=75\ \dfrac{m}{s}" is the final velocity of the fighter jet, "a" is the acceleration of the fighter jet and "s=100\ m" is the length of the aircraft carrier deck.

Then, from this equation we can find the acceleration of the fighter jet:


"a=\dfrac{v^2}{2s}=\dfrac{(75\ \dfrac{m}{s})^2}{2\cdot 100\ m}=28.1\ \dfrac{m}{s^2}."

Answer:

"a=28.1\ \dfrac{m}{s^2}."